# Robust Semiparametric Graphical Models with Skew-Elliptical Distributions

### v2

@@ 4.1 The Dantzig selector
The Dantzig selector estimator (Yuan,, 2010) exploits the relation between the coefficients $\theta$ stemming from a multivariate linear regression and the elements of the precision matrix. A sparse precision matrix implies sparse regression coefficients estimation that relies on an algorithm that depends on the tuning parameter $\delta$. The Dantzig procedure involves two steps:
@@
    $\hat{\Omega}^{gDS}=\text{arg}\min_{\Omega=\Omega^{T}}\|\Omega-\hat{\Omega}\|_{1}.$ (10)
@@ 4.2 The Glasso Algorithm
Glasso algorithm is developed by Friedman et al., (2008) and it consists in adding a $L_{1}$-norm penalty term in the maximum likelihood estimation of $\Omega$ as follows:
@@
The penalty term, $\lambda_{G}$, is the parameter that controls the sparsity of the estimated matrix, i.e., the number of zeros in the precision matrix. Banerjee et al., (2008) shows that the shrinking procedure based on the tuning parameter $\lambda_{G}$ is directly proportional to the input covariance matrix. This means that greater values of $\lambda_{G}$ imply a greater presence of zeros in the estimated $\Omega$, therefore, a minor number of interconnections are unveiled.
@@ 4.3 The CLIME Algorithm
CLIME stands for constrained $\ell_{1}$-minimization for inverse matrix estimation. This algorithm, proposed by Cai et al., (2011), is very attractive for high-dimensional data since the estimator is obtained by combining vector solutions by solving a linear program. Let $\Omega$ denote a $p\times p$ precision matrix and $\{\hat{\Omega}_{1}\}$ be the solution set of the optimization problem:
@@
where $\lambda_{C}$ is the tuning parameter. The final CLIME estimator of $\Omega$ is defined by the symmetrization of $\hat{\Omega}_{1}$:
$\hat{\Omega}=(\hat{\omega}_{ij}),\quad\text{where:}$ (13)
$\hat{\omega}_{ij}=\hat{\omega}_{ji}=\hat{\omega}^{1}_{ij}I\{|\hat{\omega}^{1}_{ij}\leq|\hat{\omega}^{1}_{ij}|\}+\hat{\omega}^{1}_{ji}I\{|\hat{\omega}^{1}_{ji}\leq|\hat{\omega}^{1}_{ji}|\}.$
@@
where $\boldsymbol{e}_{i}$ is defined as a standard unit vector in $\mathbb{R}^{p}$ with 1 in the $i$-th coordinate and 0 in the all other coordinates and $\boldsymbol{\beta}$ is a vector in $\mathbb{R}^{p}$. Solving (14) $p$ times, we obtain the CLIME estimator: $\{\hat{\Omega}_{1}\}=\{\hat{B}\}:=\{(\hat{\boldsymbol{\beta}}_{1},...,\hat{\boldsymbol{\beta}}_{p})\}$.
@@ 5 Theoretical properties
We investigate the theoretical properties of the elliptical skew-(S)KEPTIC estimators, establishing concentration inequalities in $\|\cdot\|_{\max}$ for our correlation matrix estimators, and the sparsistency of $\hat{\Omega}$ which describes the characteristic where any parameter that is originally zero is correctly estimated as zero, with a probability approaching one. For the rate of convergence of the correlation matrix estimator, we revisit the results of Liu et al., 2012a . For the corresponding precision matrix, we follow the framework introduced in Lam and Fan, (2009) for demonstrating the rate of convergence and the sparsistency of $\hat{\Omega}$. Our findings are based on the application of shrinkage procedures incorporating an $L_{1}$-penalty, with the three methods outlined in Section 4 adhering to this approach.
@@
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha},h^{(2p)},0,D_{\kappa};f)$, restricted to the case $h^{(p)}=\phi^{(p)}(u)$, and let $\hat{S}^{\rho,\alpha}_{ij}=4\sin(\frac{\pi}{6}A_{\hat{\rho,}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-SKEPTIC, then with probability at least $1-\frac{1}{p}$ we have:
$\|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq\sqrt{\frac{3(2\pi-3)^{2}}{(9-2\pi)^{2}}\frac{\log\left({p}\right)}{n}}.$ (15)
@@
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha},h^{(2p)},0,D_{\kappa};f)$ and let $\hat{S}^{\tau,\alpha}_{ij}=2\sin(\frac{\pi}{2}A_{\hat{\tau}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-(S)KEPTIC estimators, then with probability at least $1-\frac{1}{p}$ we have:
$\|\hat{S}^{\tau,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq\sqrt{{6\pi^{2}}\frac{\log\left({p}\right)}{n}}.$ (16)
The proofs of Theorems 5.1 and 5.2 are in Appendix A. The elliptical skew-(S)KEPTIC estimators are consistent, so when the sample size $n$ increases, the estimator converges to the true correlation matrix.
Let $\mathcal{S}=\{(i,j):\omega^{0}_{ij}\neq 0\}$ be the support set of off-diagonal nonzero entries, with cardinality $|\mathcal{S}|=s+p$, let $p_{\lambda}(\cdot)$ be a penalty function, and define
$a=\max_{(i,j)\in\mathcal{S}}p^{\prime}_{\lambda}(|\ell_{ij}|),\quad b=\max_{(i,j)\in\mathcal{S}}p^{\prime\prime}_{\lambda}(|\ell_{ij}|).$
where $\ell_{ij}$ is the $(i,j)$ element of the $L$ matrix that stems from $\Omega=L^{\top}ML$. In this work, we consider shrinkage procedures based on $L_{1}$-penalty, which means $a=\lambda$ and $b=0$. We impose the following regularity conditions:
  • (a)
    The eigenvalues of $\Sigma_{0}$ are bounded: $0<\tau_{1}<\lambda_{\min}(\Sigma_{0})\leq\lambda_{\max}(\Sigma_{0})<\tau_{2}<\infty$.
  • (b)
    Assume $a=O\left(\sqrt{\frac{(1+p/(s+1))\log p}{n}}\right)$, $b=o(1)$, and $\min_{(i,j)\in\mathcal{S}}|\omega^{0}_{ij}|/\lambda\to\infty$.
  • (c)
    The penalty is singular at the origin: $\lim_{t\downarrow 0}p_{\lambda}(t)/(\lambda t)=k>0$.
  • (d)
    For some constants $C,D>0$, when $\theta_{1},\theta_{2}<C\lambda$,
    $|p^{\prime\prime}_{\lambda}(\theta_{1})-p^{\prime\prime}_{\lambda}(\theta_{2})|\leq D|\theta_{1}-\theta_{2}|.$
@@
Suppose that conditions (a), (b), (c), (d) hold. If $\frac{\log p}{n}=O(\lambda^{2})$, then there exists a local minimizer for the maximum likelihood estimator for $\widehat{{M}}$ and an estimate of $\widehat{{L}}$ that satisfy:
$\|\widehat{{M}}-{M}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),\quad\|\widehat{{L}}-{L}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),$
which implies:
$\|{\Omega}-\Omega\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right).$
The proof of Theorem 5.3 is in Appendix A. We know from Lemma 1 in Zareifard et al., (2016) that $L_{ij}=0\Leftrightarrow\Omega_{ij}=0$, so the matrices $L$ and $\Omega$ share the same sparsity patter. For this reason, we provide sparsistency directly on $\Omega$.
@@
(Sparsistency) Under the conditions (a), (b), (c), (d) for a local minimizer based on $L_{1}$-penalty that satisfies $\|\widehat{\Omega}-\Omega\|_{\max}^{2}=O_{P}\left\{s\log p/n\right\}$ and $\|\widehat{\Omega}-\Omega\|^{2}=O_{P}\left(\eta_{n}\right)$ for a sequence of $\eta_{n}\rightarrow 0$, where $\eta_{n}\equiv s\sqrt{\frac{\log p}{n}}$. If $\log p/n+\eta_{n}=O\left(\lambda_{n1}^{2}\right)$, then with probability tending to $1,\widehat{\omega}_{ij}=0$ for all $(i,j)\in S^{c}$.
The proof of Theorem 5.4 is in Appendix A.
@@ 6.1 Numerical Simulations
We perform numerical simulations by generating synthetic data from the normal distribution, the closed skew-normal distribution, and the unified skew-$t$ distribution. For each dataset, we compute both the SKEPTIC and the skew-(S)KEPTIC estimators, and we evaluate the graph recovery performance of the proposed methods.
@@
The implementation follows these steps. Using the R package huge (Jiang et al., (2021)), we generate the precision matrix $\Omega=(\Sigma^{0})^{-1}$ using the ”random” structure, in which we randomly generate pairs of conditional dependencies in the off-diagonal elements. We sample $n$ observations from $p$-variate normal, CSN and SUT distribution. We set the location parameter equal to zero, and $\Sigma^{0}=\Omega^{-1}$.
We consider the following cases: $p\in\{100,250,300\}$, and $n\in\{200,1000,500\}$. We generate synthetic data assuming $f(\boldsymbol{X})\sim N(0,\Sigma^{0})$, $f(\boldsymbol{X})\sim\operatorname{CSN}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})$, and $f(\boldsymbol{X})\sim\operatorname{SUT}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})\equiv\frac{W}{\sqrt{V_{0}}}$, where $W$ is distributed according to a closed skew-normal distribution, and $V_{0}\sim Gamma(\nu/2,\nu/2)$ as in Proposition 2.5. We set $\nu=3$ to emphasize the heavy tails in the distribution. To generate the data from the closed skew-normal distribution, we rely on the stochastic representation in (7). Moreover, we generate the skewness parameters from $\alpha=0.4\cdot\operatorname{sign}(u-0.5)\cdot\sqrt{|u-1|},\quad\text{where }u\sim\mathcal{U}(0,1)$. These parameters are encoded in the diagonal of the matrix $D_{\alpha}$.
@@
After the data generation, we compute the skew-(S)KEPTIC and the SKEPTIC estimators, which are then plugged into the glasso algorithm in (11) to obtain $\Omega$. The glasso algorithm depends on a tuning parameter $\lambda_{G}$ whose the choice is performed applying the StARS method in Liu et al., (2010), and when the value is found for each dataset we consider a set of 30 tuning parameters in the range $(\lambda_{G}-0.1,\lambda_{G}+0.1)$.
To evaluate the accuracy of the methods, we use the receiver operating curve ($\operatorname{ROC}$). We define the false positive rate and true positive rate for a fixed set of tuning parameters $\lambda_{G}$. When the tuning parameter is determined we estimate the graph: $\hat{G}^{\lambda_{G}}=(V,\hat{E}^{\lambda_{G}})$, to evaluate it, we define the number of false positives for a given $\lambda_{G}$, $\operatorname{FP}(\lambda_{G})$, as the number of edges in $\hat{E}^{\lambda_{G}}$ but not in $E$ and the number of false negatives, $\operatorname{FN}(\lambda_{G})$, as the number of edges in $E$ but not in $\hat{E}^{\lambda_{G}}$. We consider the rate of the number of false negatives and false positives:
@@
Graphically, we show averaged ROC curves for all the considered scenarios over 100 trials, plotted using the coordinates in (18). Fig. 1 shows that the skew-(S)KEPTIC estimators generally outperform the Pearson correlation approach in non-Gaussian and improve upon the standard SKEPTIC estimator. Adjusting for skewness, also for excess kurtosis in the case of the SUT distribution, enhances the estimation of $\Omega$, especially in settings where the distribution presents heavy tails. In some cases, this leads to better recovery compared to the original SKEPTIC estimator, which has been shown to be optimal in Liu et al., 2012a . Table 2 provides a comprehensive overview of the simulation results, presenting quantitative comparisons of various estimators. The average AUC values of the skew-(S)KEPTIC estimators highlight strong performance in graph recovery under non-Gaussian conditions while still maintaining reliability when applied to normally distributed data. These findings suggest that considering estimators that include the skewness and accommodate for heavier tails, extending the nonparanormal distribution, do not hinder the recovery of the $\Omega$ structure. On the contrary, they may offer a practical advantage due to the robustness across different data distributions.
@@
Figure 1: Comparison of average ROC curves, using the glasso algorithm, of the elliptical skew-(S)KEPTIC estimators with SKEPTIC estimators. Results are determined over 50 trials. We consider, by row, three different data generations and, by column, three different dataset sizes.

### v3

@@ 4.1 The Dantzig selector
The Dantzig selector estimator (yuan2010high) exploits the relation between the coefficients $\theta$ stemming from a multivariate linear regression and the elements of the precision matrix. A sparse precision matrix implies sparse regression coefficients estimation that relies on an algorithm that depends on the tuning parameter $\delta$. The Dantzig procedure involves two steps:
@@
    $\widehat{\Omega}^{gDS}=\text{arg}\min_{\Omega=\Omega^{T}}\|\Omega-\widehat{\Omega}\|_{1}.$ (10)
@@ 4.2 The Glasso Algorithm
Glasso algorithm is developed by friedman2008sparse and it consists in adding a $L_{1}$-norm penalty term in the maximum likelihood estimation of $\Omega$ as follows:
@@
The penalty term, $\lambda_{G}$, is the parameter that controls the sparsity of the estimated matrix, i.e., the number of zeros in the precision matrix. banerjee2008model shows that the shrinking procedure based on the tuning parameter $\lambda_{G}$ is directly proportional to the input covariance matrix. This means that greater values of $\lambda_{G}$ imply a greater presence of zeros in the estimated $\Omega$, therefore, a minor number of interconnections are unveiled.
@@ 4.3 The CLIME Algorithm
CLIME stands for constrained $\ell_{1}$-minimization for inverse matrix estimation. This algorithm, proposed by cai2011constrained, is very attractive for high-dimensional data since the estimator is obtained by combining vector solutions by solving a linear program. Let $\Omega$ denote a $p\times p$ precision matrix and $\{\widehat{\Omega}_{1}\}$ be the solution set of the optimization problem:
@@
where $\lambda_{C}$ is the tuning parameter. The final CLIME estimator of $\Omega$ is defined by the symmetrization of $\widehat{\Omega}_{1}$:
$\widehat{\Omega}=(\widehat{\omega}_{ij}),\quad\text{where:}$ (13)
$\widehat{\omega}_{ij}=\widehat{\omega}_{ji}=\widehat{\omega}^{1}_{ij}I\{|\widehat{\omega}^{1}_{ij}\leq|\widehat{\omega}^{1}_{ij}|\}+\widehat{\omega}^{1}_{ji}I\{|\widehat{\omega}^{1}_{ji}\leq|\widehat{\omega}^{1}_{ji}|\}.$
@@
where $\boldsymbol{e}_{i}$ is defined as a standard unit vector in $\mathbb{R}^{p}$ with 1 in the $i$-th coordinate and 0 in the all other coordinates and $\boldsymbol{\beta}$ is a vector in $\mathbb{R}^{p}$. Solving (14) $p$ times, we obtain the CLIME estimator: $\{\widehat{\Omega}_{1}\}=\{\widehat{B}\}:=\{(\widehat{\boldsymbol{\beta}}_{1},...,\widehat{\boldsymbol{\beta}}_{p})\}$.
@@ 5 Theoretical properties
We analyze the theoretical properties of the elliptical skew-(S)KEPTIC estimator. The results suggest that our estimator is reliable when dealing with highly skewed data. Our results show an exponential concentration rate in $\|\cdot\|_{\max}$. Moreover, we prove optimal bounds for shrinkage procedures. We show the results for the Dantzig selector, however, they can be easily extended to the CLIME and the Glasso procedures.
@@
Let $x_{1},\dots,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},\Lambda,h^{(p)},0,\Gamma;f)$ restricted to the case $h^{(p)}=\phi^{(p)}(u)$, and let
$\widehat{S}^{\rho,\alpha}_{ij}=2\sin\!\left(\tfrac{\pi}{6}\widehat{\rho}_{ij}\right)\,\frac{1+\widehat{\alpha}_{i}\widehat{\alpha}_{j}}{\sqrt{(1+\widehat{\alpha}_{i}^{2})(1+\widehat{\alpha}_{j}^{2})}}\,\operatorname{I}(i\neq j)\;+\;\operatorname{I}(i=j)$
denote the elliptical skew-SKEPTIC estimator. Then, with probability at least $1-2/p^{2}$, we have
$\|\widehat{S}^{\rho,\alpha}-\Sigma^{0}\|_{\max}\leq(8\pi+2C_{\alpha})\sqrt{\frac{\log p}{n}}.$ (15)
@@
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},\Lambda,h^{(p)},0,\Gamma;f)$ and let
$\widehat{S}^{\tau,\alpha}_{ij}=\sin\!\left(\tfrac{\pi}{2}\widehat{\tau}_{ij}\right)\,\frac{1+\widehat{\alpha}_{i}\widehat{\alpha}_{j}}{\sqrt{(1+\widehat{\alpha}_{i}^{2})(1+\widehat{\alpha}_{j}^{2})}}\,\operatorname{I}(i\neq j)\;+\;\operatorname{I}(i=j),$
denote the elliptical skew-(S)KEPTIC estimator, then with probability at least $1-2/p^{2}$ we have:
$\|\widehat{S}^{\tau,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq(2.45\pi+2C_{\alpha})\sqrt{\frac{\log p}{n}}.$ (16)
Proofs of Theorems 5.1 and 5.2 are in A. The elliptical skew-(S)KEPTIC estimator achieves convergence, thus it is possible to define the following result.
@@
When we substitute the estimated matrices $\hat{S}^{\rho,\alpha}$ and $\hat{S}^{\tau,\alpha}$ into the parametric Graphical Lasso (or the CLIME or graphical Dantzig selector), the estimators provide convergence, ensuring consistency and graph recovery.
Proof.
The proof is based on the observation that the sample correlation matrix $\widehat{S}$ is a sufficient statistic for all three shrinkage procedures: the graphical lasso, graphical Dantzig selector, and CLIME. As yuan2010high points out, a sufficient condition to ensure that an estimator is consistent is if there exists some constant $c$, such that $\mathbb{P}\left(\|\widehat{S}-\Sigma^{0}\|_{\max}>c\sqrt{\frac{\log p}{n}}\right)\leq 1-\frac{1}{p}$, which can be replaced by (15) and (16) in Theorems 5.1 and 5.2. ∎
The convergence of the elliptical skew-(S)KEPTIC estimator also means that the precision matrix is bounded. Similar results are found by cai2011constrained and yuan2010high. In particular, let $\widehat{\Omega}^{mse-s}$ denote the inverse correlation matrix using the elliptical skew-(S)KEPTIC estimator and consider the Dantzig selector as the shrinkage procedure to find $\widehat{\Omega}$. We introduce a class of inverse matrix defined as $\mathcal{M}_{1}(\kappa,\tau,M):=\{\Omega:\Omega\succ 0,\text{diag}(\Omega^{-1})=1,\|\Omega\|_{1}\leq\kappa,\frac{1}{\tau}\leq\lambda_{\min}(\Omega)\leq\lambda_{\max}(\Omega)\leq\tau,\deg(\Omega)\leq M\}$, where $\kappa,\tau>1$. Given Theorem 5.3, we obtain the following result:
@@
For $1\leq q<\infty$, there exists a constant $c$ that depends on $\kappa$, $\tau$, $\lambda_{\min}(\Omega)$ and $\lambda_{\max}(\Omega)$, such that
$\sup_{\Omega\in\mathcal{M}_{1}(\kappa,\tau,M)}\|\widehat{\Omega}^{mse-s}-\Omega\|_{q}=O_{P}\left(M\sqrt{\frac{\log p}{n}}\right).$ (17)
provided that $\lim_{n\rightarrow\infty}\frac{n}{M^{2}\log d}=\infty$ and $\delta=c\sqrt{\frac{\log d}{n}}$, for sufficiently large $c$, where $\delta$ is the parameter in (9).
Proof.
This result follows yuan2010high, which proves that the difference between the estimated precision matrix and the theoretical precision matrix is bounded by $\|\widehat{S}-\Sigma^{0}\|_{\max}$, where $\widehat{S}$ is the estimated correlation matrix and $\Sigma^{0}$ is the theoretical correlation matrix. In our case, to prove the bounds for the estimated correlation matrix, we just replace $\widehat{S}$ with $\hat{S}^{\rho,\alpha}$ and $\hat{S}^{\tau,\alpha}$ where the bounds are proved in Theorem 5.1 and 5.2. ∎
@@ 6 Experiments
In this section, we conduct an empirical analysis based on both synthetic and real data. The study of synthetic data is performed through numerical simulations, whereas the analysis of real data involves the estimation of undirected graphs for the stocks that compose the S&P500500 index.
@@ 6.1 Numerical Simulations
The implementation of the numerical simulations is carried out following these steps. Using the R package huge (jiang2020package), we generate the precision matrix $\Omega=(\Sigma^{0})^{-1}$ using the ”random” structure, in which we randomly generate pairs of conditional dependencies in the off-diagonal elements, the sparsity of $\Omega$ is set to 0.02 to ensure the positive defitivness of the matrix so that the matrix can be inverted. We consider the case of $p=100$, and $n=200$, and we sample $n$ observations from $p$-variate normal distribution. We set the location parameter equal to zero, and $\Sigma^{0}=\Omega^{-1}$.
To evaluate the robustness of our estimators, we introduce a random contamination mechanism. For each marginal distribution, let $r\in(0,1)$ denote the contamination level. We randomly select $\lfloor nr\rfloor$ observations and replace them with values drawn from $\mathcal{Y}\sim N(0,5)$. Moreover, we apply the power transformation to the contaminated data, which is the following. Let $g_{0}(t):=\operatorname{sign}(t)|t|^{\gamma}$ where $\gamma>0$ is a parameter. The power transformation for the $j$th dimension is defined as
$g_{j}^{0}\left(z_{j}\right):=\frac{g_{0}\left(z_{j}-\mu_{j}\right)}{\sqrt{\int g_{0}^{2}\left(t-\mu_{j}\right)\phi\left(\left(t-\mu_{j}\right)/\sigma_{j}\right)dt}},$
where $\phi(\cdot)$ is the standard Gaussian density function. These two procedures applied to the data induce a departure from normality, resulting in marginal distributions characterized by asymmetry and heavier tails compared to the Gaussian case. We consider three contamination levels (r = 0.05, 0.1, and 0.2), and three different parameters of the power function ($\gamma$ = 1.5, 2.5, and 3), considering a total of nine scenarios.
In this analysis, we want to estimate the asymmetry of the data estimating the skewness of the data encoded in the matrix $D_{\alpha}$ of the distributions $f(\boldsymbol{X})\sim\operatorname{CSN}_{p,p}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})$, and $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})$. Regarding the closed skew-normal, we use the EM algorithm described in abe2021algorithm, for the unified skew-$t$ we employ the likelihood-based method from the R package sn developed by azzalini_pkg_sn.
@@
After the data generation, we compute the skew-(S)KEPTIC and the SKEPTIC estimators, which are then plugged into the glasso algorithm in (11) to obtain $\Omega$. The glasso algorithm depends on a tuning parameter $\lambda_{G}$ whose choice is performed by applying the StARS method in liu2010stability, which selects the regularization parameter based on the stability of the estimated graph under subsampling, and when the value is found for each dataset, we consider a set of 30 tuning parameters in the range $(\lambda_{G}-0.1,\lambda_{G}+0.1)$. To evaluate the accuracy of the methods, we use the receiver operating curve ($\operatorname{ROC}$). We define the false positive rate and true positive rate for a fixed set of tuning parameters $\lambda_{G}$. When the tuning parameter is determined we estimate the graph: $\widehat{G}^{\lambda_{G}}=(V,\widehat{E}^{\lambda_{G}})$, to evaluate it, we define the number of false positives for a given $\lambda_{G}$, $\operatorname{FP}(\lambda_{G})$, as the number of edges in $\widehat{E}^{\lambda_{G}}$ but not in $E$ and the number of false negatives, $\operatorname{FN}(\lambda_{G})$, as the number of edges in $E$ but not in $\widehat{E}^{\lambda_{G}}$. We consider the rate of the number of false negatives and false positives:
@@
Graphically, we show averaged ROC curves for all the considered scenarios over 50 trials, plotted using the coordinates in (19). Fig. 1 illustrates that the skew-(S)KEPTIC estimators yield results comparable to those of the SKEPTIC estimator, which has been shown in liu2012high to be optimal for a similar class of data. Importantly, incorporating skewness does not deteriorate the recovery of the true precision matrix $\Omega$. A more detailed comparison is provided in Table 2, which reports the values (in percentage) of the Area Under the Curve (AUC), False Positive Rate (FPR), and False Negative Rate (FNR). Across all scenarios, taking into account both data contamination and power-function transformations, the elliptical skew-(S)KEPTIC estimators exhibit lower false positive rates and higher false negative rates, reflecting the smaller off-diagonal entries relative to those of the SKEPTIC estimator. Given that the tuning parameter $\lambda_{G}$ is fixed across all methods, it is not expected that the elliptical skew-(S)KEPTIC estimators outperform the SKEPTIC estimator in terms of overall graph recovery. Nevertheless, it is noteworthy that the skew-(S)KEPTIC estimators consistently exhibit superior performance compared to the Pearson correlation coefficient in the recovery of graphical structures.
These findings suggest that considering estimators that include the skewness of a skew-normal or a skew-t distribution and extend the nonparanormal SKEPTIC estimator, do not hinder the recovery of the $\Omega$ structure. On the contrary, they may offer a practical advantage due to the robustness across different data distributions.
@@
Figure 1: Comparison of average ROC curves, using the glasso algorithm, of the elliptical skew-(S)KEPTIC estimators with the SKEPTIC estimator. Results are determined over 50 trials. We consider, by row, three different data contamination and, by column, three different values of the parameter $\gamma$ that drives the power transformation.